OpenStudy (anonymous):

can someone please show me step by step how to get the point slope and slope intercept form for passing through (-2,3) and parallel to the line whose equation is 2x-3y=5?

OpenStudy (anonymous):

i got the point slope form just not the slope intercept form cause for that i got y=2/3x+11/3 but it should be 13/3

OpenStudy (anonymous):

and i don't know how they got it

OpenStudy (anonymous):

"can someone please show me step by step how to get the point slope and slope intercept form for passing through (-2,3) and parallel to the line whose equation is 2x-3y=5?" Parallel to 2x-3y=5 means it has the same slope, viz. 2/3. That you got. The point-slope form is y-3=2/3(x-(-2)). Simplify that and move the 3 to the other side.

OpenStudy (anonymous):

Your constant will be 4/3 + 3 = 4/3 + 9/3.

OpenStudy (anonymous):

You might not have distributed the 2/3 to the 2 properly . . .

OpenStudy (anonymous):

oh okay